OpenStudy (yadu123):

a. I have added 15,000mL of air to a balloon at sea level. If I travel to Flagstaff, Arizona, where the air pressure is 0.85atm, what will the balloon’s new volume be?

OpenStudy (welshfella):

P1*V1 = P2*V2 where P1 and V1 are volume and pressure at sea level and P2V2 are the values in arizona

OpenStudy (welshfella):

we have 15,000 * 1 = V2 * 0.85 solve for V2

OpenStudy (yadu123):

ok makes sense

OpenStudy (welshfella):

yes I think its called Boyles Law for gases

OpenStudy (yadu123):

17,647 mL

OpenStudy (yadu123):

Thx

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

yw